A certain organ pipe, open at both ends, produces a fundamental frequency of 271 Hz in air.
If the pipe is filled with helium at the same temperature, what fundamental frequency f He will it produce? Take the molar mass of air to be 28.8 g/mol and the molar mass of helium to be 4.00g/mol .
Answer #1
Concepts and reason

The concepts required to solve this problem are the velocity of waves in a medium and the relation between the velocity and frequency.

Use the expression for velocity of the wave in a medium and the velocity and frequency relation to calculate the fundamental frequency in the Helium medium.

Fundamentals

Write the expression for the velocity of the wave in a medium.

c=γRTMc = \sqrt {\frac{{\gamma RT}}{M}}

Here, RR is the gas constant, γ\gamma is the adiabatic index, TT is the absolute temperature and MM is the Molar mass.

Write the expression for the frequency of the wave:

f=cλf = \frac{c}{\lambda }

Here, λ\lambda is the wavelength and cc is the speed of the wave.

The velocity of the wave in a medium is given by the relation as follows:

c=γRTMc = \sqrt {\frac{{\gamma RT}}{M}}

Here, RR is the gas constant, γ\gamma is the adiabatic index, TT is the absolute temperature and MM is the Molar mass.

Write the expression for the frequency of the wave:

f=cλf = \frac{c}{\lambda }

Here, λ\lambda is the wavelength and cc is the speed of the wave.

Therefore, by the expression of frequency and velocity it can be deduced that the frequency of the wave is given by the relation as follows:

f=1λγRTMf = \frac{1}{\lambda }\sqrt {\frac{{\gamma RT}}{M}}

The obtained relation of frequency of the wave is given by the relation as follows:

f=1λγRTMf = \frac{1}{\lambda }\sqrt {\frac{{\gamma RT}}{M}}

Here, λ\lambda is the wavelength, RR is the gas constant, γ\gamma is the adiabatic index, TT is the absolute temperature and MM is the molar mass.

For air, the equation for the frequency of the wave can be written as follows:

fair=1λγairRTMair{f_{{\rm{air}}}} = \frac{1}{\lambda }\sqrt {\frac{{{\gamma _{{\rm{air}}}}RT}}{{{M_{{\rm{air}}}}}}}

Here, γair{\gamma _{{\rm{air}}}} is the adiabatic index of the air and Mair{M_{{\rm{air}}}} is the molar mass of the air.

For helium, the expression for the frequency in the helium medium can be written as follows:

fHe=1λγHeRTMHe{f_{{\rm{He}}}} = \frac{1}{\lambda }\sqrt {\frac{{{\gamma _{{\rm{He}}}}RT}}{{{M_{{\rm{He}}}}}}}

Here, γHe{\gamma _{{\rm{He}}}} is the adiabatic index of Helium and MHe{M_{{\rm{He}}}} is the molar mass of Helium.

Take the ratio of the frequency in helium to the frequency in air.

fHefair=1λγHeRTMHe1λγairRTMair\frac{{{f_{{\rm{He}}}}}}{{{f_{{\rm{air}}}}}} = \frac{{\frac{1}{\lambda }\sqrt {\frac{{{\gamma _{{\rm{He}}}}RT}}{{{M_{{\rm{He}}}}}}} }}{{\frac{1}{\lambda }\sqrt {\frac{{{\gamma _{{\rm{air}}}}RT}}{{{M_{{\rm{air}}}}}}} }}

Rearrange for fHe{f_{{\rm{He}}}} .

fHe=fairγHeMairγairMHe{f_{He}} = {f_{{\rm{air}}}}\sqrt {\frac{{{\gamma _{{\rm{He}}}}{M_{{\rm{air}}}}}}{{{\gamma _{{\rm{air}}}}{M_{{\rm{He}}}}}}}

Substitute 271Hz271{\rm{ Hz}} for fair{f_{{\rm{air}}}} , 1.41.4 for γair{\gamma _{{\rm{air}}}} , 28.8g/mol28.8{\rm{ g/mol}} for Mair{M_{{\rm{air}}}} , 1.671.67 for γHe{\gamma _{{\rm{He}}}} and 4g/mol4{\rm{ g/mol}} for MHe{M_{{\rm{He}}}} .

fHe=(271Hz)(1.67)(28.8g/mol)(1.4)(4g/mol)=(271Hz)8.59=(271Hz)(2.93)=794.2Hz\begin{array}{c}\\{f_{{\rm{He}}}} = \left( {271{\rm{ Hz}}} \right)\sqrt {\frac{{\left( {1.67} \right)\left( {28.8{\rm{ g/mol}}} \right)}}{{\left( {1.4} \right)\left( {4{\rm{ g/mol}}} \right)}}} \\\\ = \left( {271{\rm{ Hz}}} \right)\sqrt {8.59} \\\\ = \left( {271{\rm{ Hz}}} \right)\left( {2.93} \right)\\\\ = 794.2{\rm{ Hz}}\\\end{array}

Ans:

The fundamental frequency produced by the organ pipe in Helium medium is 794.2Hz{\rm{794}}{\rm{.2 Hz}} .